OpenStudy (dller008):

Question 5 (Multiple Choice Worth 2 points) (09.03) Jean bought a bag of candy that contained red candies, green candies, and brown candies. Some candies are marked with a smile, and others are not marked. The probabilities are displayed in the table below: Red Green Brown Smile 0.10 0.30 0.10 No smile 0.20 0.15 0.15 If a random candy is chosen from this bag, what is the probability that the candy was red or had no smile? 0.60 0.20 0.30 0.50

OpenStudy (gunsodo):

Hint Red or had no smile = 0.10+0.20+0.15+0.15

OpenStudy (wanell):

| Red |Green | Brown Total Smile |0.100 | .300 | .10 0.50 No smile |0.200 | .150 | .15 0.50 Total 0.30 0.45 0.25 1.00 The probability of choosing red candy is 0.10 The probability of choosing no smile candy is 0.50 The probability of choosing red or no smile = 0.10 + 0.50 = 0.60 I did not use 0.30 as the probability of red because it included the no smile candy. It should only be red candy that has a smile.